19 September 2010

Afternoon:
  • Assessed focal plane tilt. I suspect a small ~50 micron, tilt in Blue, such that chip 1 is closer to M1 than chip 3. But need to check data more carefully.
  • Obtained new collimation models for Blue and Red. Data are in /home/kuhn/LBTO/20100918/CollimData and in /home/lbcobs/Collimation/new and the new models are called SXPMLBCCollimation20100918.dat and DXPMLBCCollimation20100918.dat. I copied these to /home/telescope/TCS/Configuration/new/ and replaced the existing defaults by these.
  • Measured rotator center from one pair of PA=0/PA=180 images. The Blue rotator center changed by (+6,-3) and the Red rotator center changed by (-3,-3) (=below are not the final rotator centers that I entered in the headers - see below).
    • Old Red crpix1,crpix2: New Red crpix1,crpix2
         1: -1042 2913  -1045 2910
         2: 1080 2913   1077 2910
         3: 3202 2913  3199 2910
         4: -1719 2237 -1722 2240
    • Old Blue crpix1, crpix2 New Blue crpix1, crpix2
         1: -1090 2921 -1084 2919
         2: 1032 2921 1038 2919
         3: 3154 2921  3160 2919
         4: -1711 2274 -1714 2268
  • 02:04 Try out myskyflat.pl script (in /home/LBTO/Calib_OBs) which will generate a flat field OB. Works fine.
  • 02:21 Collimating on first field. Now the first Red pupils are fit by FPIA. No manual corrections are needed. (Yesterday we needed to give Z7=Z8=1500). The red pupils are still not ideal, there is a lot of astigmatism. T=13.8, so about 1.8 C warmer than when we obtained the collimation model data.
  • Took about 11 minutes to converge ---estimated seeing is about 2".
  • 023449/023446 B/R PA=0 5sec images to check seeing and rot center again.
  • 023739/023733 B/R PA=180 B rotcenter from these (1038.7,2920) R rotcenter from these (1078,2910). Thse are using averages, what I did in the afternoon bc geomap did not give reasonable results. But now geomap does. Results are:
     for Blue:
      Xc=1042.49 Yc=2918.92 from geomap.
      Xc=1038.64 Yc=2920. from averages.
      for Red:
      Xc=1081.07 Yc=2909.22 from geomap.
      Xc=1078.12 Yc=2910.27 from averages.
Keep Rotator Center as is in the header.
  • Actually seeing is very good: B 3.1 pix and R 3.3 pix --- R (r-SLOAN) has obvious IQ problem - trefoil, due to bad actuator? pr just poor collimation. FPIA was estimating seeing of about 2" - usually much closer to actual seeing than that.
  • to WT10_302 - collimating. Preparing to take pointing model data. Air and steel within 0.5 deg. Pretty close to thermal equilibrium.
  • Again - check rotator center. PA=0 030930/030927 and for PA=180 031221/031217
    • this time the rotator centers for both blue and red were different from what I measured yesterday and earlier tonight, by -3,-5 (B) and -4,-7 (R). Geomap and Avg were different as before, but close (not unreasonable). More stars than before - bright pointing star.
  • Correct pointing using Red PA=0 image 030927. Elev>80 deg. delta IE CA are 3, -31" and NEW CA, IE: -30.9 -34.2.
  • Take 2 1sec images here at PA=0 and 2 with PA=180. PA = 0: 032641/032642 and 032806/032807 PA=180: 033031/033027 and 033128/033129.
R 032807 and 033129: geomap std~0.1 and results are
      Xc=1080.14 Yc=2911.47 from geomap.
      Xc=1079.66 Yc=2911.72 from averages.
B 032642 and 032806: geomap std~0.2,0.3      
      Xc=1040.18 Yc=2921.3 from geomap.
      Xc=1039.17 Yc=2921.49 from averages.
  • Adopt RC(Red) = 1080,2912 and RC(Blue) = 1040,2921
  • Enter these values and all chip RPIX in headers on CMU.
  • LBC does not need to be restarted - these values are in headers now.
  • Slew to ACT0304 to check/correct pointing and co-pointing at high elevation, prior to collecting pointing model data. Elev=84.7deg.
  • Collimating... called Dave to check what sort of initial coll/co-pointing we need and whether to collimate between stars. We will collimate and co-point to start (on ACT0304) but thereafter will not collimate - just point and shoot.
  • 041124/041121 are B/R ACT0304. mirradjust on these ===>
dX = -0.87 mm, dY = -0.22 mm, dRX = -4.73 arcsec and dRY = 18.20 arcsec (remember tip = dRX and tilt = -dRY --- Steve is inputting tip = -5" and tilt= -18".)
  • new pair of images: 041558/041554 --- Looks good.
  • Starting to Collect Pointing Model Data
    • Steve starting log on TCS: 20100919.ptlog
    • WT10_273... no, should have gone to WT10_373... Slewing to WT10_373.
  • 1 WT10_373 4:23 042415/042410
  • 2 ACT0407 4:26 042732/042727
  • 3 WT10_401 4:28 042958/042953 (SX T=13.8 ambient 14.2 glass DX 13.7 ambient, 13.6 glass)
  • 4 WT10_358 4:31 043259/043256 (steel temps are about 14)
  • 5 ACT0281 4:34 043621/043616 (+zlimit hit on SX; we are at elev=19 deg)
  • 6 ACT0179 4:37 043913/043909
  • 7 WT10_253 4:40 044217/044212
  • 8 ACT0083 4:43 (hit focus limit on both SX and DX,both +ve limits. Elev = 21 deg) 044509/044505
  • 9 ACT0168 4:46 044752/044747
  • 10 WT10_274 4:48 045038/045033 outside temp LBTtemp has dropped to 12.6
    • (glass = 14.1 SX and 13.5 on DX; ambient on SX=13.4 DX=13.1 and steel 13.8/14.3 )
  • 11 ACT0208 045458/045352 elev= 21.4 deg. (hit +ve Z limit on SX and DX. SX Z=2.9 and RX/RY are red, DX Z=3.0 and Z/RX/RY are red and RX/RY=0/0 (even though this is not the sum of values).
  • 12 WT10_340 4:56 045732/045727
  • 13 WT10_338 045958/045955
  • 14 ACT0278 5:01 050229/050225
  • 15 WT10_242 5:03 050444/050440
  • 16 ACT0306 5:05 050752/050748
  • 17 ACT0185 5:09 051108/051103
  • 18 ACT0100 5:12 051340/051335 (hit limits on SX Z=2.7 and RX/RY are red - Z is not red) elev=25deg
  • 19 WT10_115 051554/051547 no limits now, elev=31.8 deg.
  • 20 ACT0289 5:16 051911/051906
  • 21 WT10_270 5:19 052146/052141
  • 22 ACT0213 5:22 052433/052428 Hit both SX and DX +ve Z limits Z/RX/RY are all red and RX/RY are zeroes for both SX/DX. Elev = 18.2 deg T=12.3 (Wow! Co-pointing way off at this low elevation no, RX/RY were not at their target positions; we are repeating this position later with a z global offset - see below.).
  • 23 WT10_300 5:26 052804/052759
  • 24 WT10_321 053127/053122
  • 25 ACT0398 5:32 053415/053410
  • 26 ACT0406 5:35 053645/053640 (Z = 2.7 but not red, RX/RY are red. RX=24 RY=-56.6 -- these are not out of range are they - are they red because Z is high? These RX/RY values represent the sums over the columns on PSF - the various contributions to RX/RY)
  • 27 WT10_410 5:39 054055/054050
  • 28 ACT0336 5:42 054436/054432
  • 29 WT10_375 5:45 054718/054713
  • 30 WT10_388 5:48 054956/054951 hit SX limit Z=2.88 and Z is not red, but RX/RY are red and = 16/-39. elev=21 deg
  • 31 WT10_400 5:50 055212/055207
  • 32 ACT0411 5:53 055450/055445
  • 33 WT10_360 5:56 055742/055737
  • 34 ACT0387 5:58 060035/060030 elev = 25deg. Same as number 30... Z is high but not red, and RX/RY are red but reasonable values.
  • 35 WT10_369 6:01 060231/060226
  • 36 WT10_358 6:03 060428/060424
  • 37 999! try again. Pointingstars5 finds no star for this cell (cell number 999)
  • 38 WT10_304 060900/060854
  • 39 ACT0217 6:10 061201/061155
  • 40 ACT0145 6:13 061435/061429 limits again, this time on both SX and DX. On DX all Z/RX/RY are red, on SX only RX/RY are red. elevation = 18.9 degrees.
  • 41 ACT0147 061656/061651
  • 42 WT10_202 061952/061947
  • 43 ACT0221 6:21 062213/062208
  • 44 WT10_204 6:22 062412/062407
  • 45 999 again - no star found
  • 46 WT10_104 062733/062729 both SX and DX limits are hit elev=19.6, +ve Z = 2.9/3 for SX/DX
  • 37 back to 37 ... WT10_321 063204/063200
  • 47 xxxx WT10_172 at elevation = 2 ?! we hit +ve Z limits of course 063444/063445 and ???
  • 47 WT10_172 ... but it is doing a long slew. Elev=35. No limits hit. This is correct on object. 064017/064012.
  • 48 WT10_261 064340/064335
  • 49 WT10_293 064600/064555
  • 50 ACT0312 064833/064829
  • 51 ACT0282 065132/065126 elevation 23 SX Z/RX/RY limits hit.
  • 52 WT10_355 065357/065353
  • 45 ACT0153 065717/065712
  • temperatures steel ~13 (differ from air by <0.5deg) SX glass 13 ambient 12.5 DX glass 12.4 ambient 12.4
  • 46repeat WT10_121 entered z global offset to SX of Z(go) = -0.2 (took data while doing this: 070024/070020 are the bogus images taken while adjusting Z(go).)
    • images after glob. off. entered are: 070128/070127 elev=23 deg
  • 40repeat WT10_164 070455/070453 elev=23 deg
  • 34repeat WT10_378 070805/070759 elev=19 deg
  • 30repeat WT10_370 --- very low... may need to adjust Z again, so for now go to a different star.
  • 26repeat ACT0406 071302/071303 elev=21 deg
  • 22repeat ACT0198 --- very low --- skip for later.
  • 18repeat WT10_074 071831/071827 elev=19.2 deg
  • 11repeat WT10_229 072118/072113 elev=25 deg
  • 8repeat WT10_072... skip
  • 5repeat WT10_285 072637/072632
  • 8yetagain WT10_072 Steve is entering a new Z(go): Z(go) on SX is now = -0.3 (from -0.2) and DX(go) = -0.2 (from 0.0)
    • WT10_072 073105/073057 elev=17.5 no limits
  • 22yetagain ACT0217 073400/073356 elev=23
  • 30yetagain ACT0370 073742/073738 elev=18.3
  • Done with taking Pointing Data

  • 07:38 On to a long R exposure to check condensation on Red window.
  • Removing the Z global offsets we had entered for these low elevation pointing stars.
  • Collimating on field. 2 Red images are missing - taken between 074628 and 075039. Not converging well. We are at low elevation and on an extragalactic field: only 1-2 pupils are fit. 9 iterations already, ~10 minutes. At 08:00, hung - no new files appearing in /newdata. Restarted dofpia. No good pupils. I offset SW. Converged in Blue after 1st try, Red was close, but I stopped with Shift-S.
  • Running OB which will take 420 sec exposure in Blue (SDT-Uspec) and Red (z-SLOAN). Saturated at red. Change to Y-FAN filter. RH=70%. 3 x 420 sec exposures. 1st one looks OK but has a strange gradient over the field. We are looking South - Moon is SW.
    • Red frame 082928 shows star trails that are 90 pixels, 20.25" long, but corresponding Blue image 082927 is OK. Red was master, Blue Slave. Elev= 28. Subsequent Red image is OK. No sign of condensation, but odd pattern.
  • Trying higher field: Little Dumbell Nebula at 70 deg elevation.
  • 08:55 Collimating. RH=74%.
  • 08:57 Run OB. Uses SDT_Uspec and Y-FAN filters. Red Master Blue Slave. FPIA estimates seeing 1.1 B and
0.9 R, but Guide thumbnails show 1.3 R 1.9 B (is this because focus offsets are wrong, or FPIA is wrong, or guide FWHM is wrong?) Actual seeing on the image is close to the guide FWHM estimate. B 6.6 (1.5"), R 5.2 (1.2").
  • There is a pupil on _2 red tech chips, but may be too close for AO? But I see in the log that some P03 corr sent to B Trackers. 09:21 and 09:13
  • First image looks OK in red - there is a similar 'scattered light' problem but much less apparent than for previous field. No sign of condensation. RH now 66.5%. * Now try same sequence with Blue master/Red slave. Blue image has FWHM~5.8 Red image has FWHM~4.8 * At 09:39 we have some P03 (Active Optics) corr sent for B and R.
    2010/09/19 09:39:33.405376 S B TRACKERS          Active Optics P03 -261.00000000 +0.00000000
    2010/09/19 09:39:34.501070 S R TRACKERS          Active Optics P03 -152.00000000 +0.00000000
    
  • I see no sign of condensation in long Red exposures. But I do see a strange scattered light pattern. This is through the Y filter. I am taking data now through r-SLOAN (and B-BESSEL). The Red images look like a bright star is just off the W of the array, but the Blue images do not show this. We'll take a look in the dome. No sign of this at r-SLOAN (RED). Trying z-SLOAN. There was a bright star off to the W.
  • Going to open star cluster NGC1039. Running dofpia. Converged in 3 iterations - 3-4 minutes.
  • Observing star cluster to watch tech chip images: Red Master, Blue Slave 101632/101627.
    • Red tech chip image 101655 has at least two bright stars, but there are no P03 corrections. (z)
    • Blue tech chip image 101656 has only 1 bright star - no P03 corrections. (U)
    • 10:23 there was a B P03 trackers correction.
  • Red RPC Client/Host communication error at 03:23. After this I just restarted the OB to see what would happen. The blue side is taking data and guide corrections continue to be sent to Z2/Z3 of primary mirror. But with red side, nothing is happening. Will turn off and turn back on LBC.
2010/09/19 10:07:27.664640 E R                   OB execution error: none [src/lbc/actions.c:1639]
2010/09/19 10:20:15.538641 E R TRACKERS          RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1074]
2010/09/19 10:21:47.191541 E R TRACKERS          reinitializing CCD Controller failed
2010/09/19 10:21:47.191619 E R TRACKERS          timeout on CCD Controller [src/trackers/trackers.c:2168]
2010/09/19 10:21:47.191696 E R                   trackers system encountered a problem
2010/09/19 10:21:47.191758 E R                   >>> preset/expose/save interrupted at expose time
2010/09/19 10:21:47.311763 E R                   OB execution error: none [src/lbc/actions.c:1639]
  • 03:30 Turning off LBC to clear RPC error on Red Trackers. Takes a while - Trackers in "error" state and
then "warning". Turning LBC back on.
  • 03:45 LBC back up, collimated. Running OB again: B-BESSEL, z-SLOAN. Both have pupils in AO chip.
  • 04:11 Launch Vgr_iYR OB on this same field to obtain tech chip pupil data at these filters. gY produce corrections for both R/B; Vi did not. R (R-B) gets correction at 11:19 and B (r-SLOAN) does too. This time Vi give corrections. At 04:28, it looks like Red focus has gone off (i-SLOAN filter 1.6"). Earlier image through i-SLOAN was
round and OK.
  • 04:35 Focus/collimate and then run one last OB - this time with Blue Master/Red Slave and with SU_IV to obtain AO chip data for these.
    • Focus/Collimate at 04:36. Previous 2 OBs took over 80 minutes. Converged in 2 iterations!
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1426    591   -428    193    146   -223      6    y  1.43  0.63                      /newdata/lbcb.20100919.113731.fits  V-BESSEL
    66   -249    337    -19   -162   -113     44    y  1.43  0.18                      /newdata/lbcb.20100919.113900.fits  V-BESSEL
   187      0      0      0      0    -98      0    y  0.00  0.00                     dofpia_sph3_adjustment             

  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   868   -201     50   -130    -84     -6      6    y  1.14  0.36                      /newdata/lbcr.20100919.113730.fits  r-SLOAN 
    95   -121    210    108    -16    -11      1    y  1.15  0.12                      /newdata/lbcr.20100919.113901.fits  r-SLOAN 
   910      0      0      0      0   -479      0    y  0.00  0.00                     dofpia_sph3_adjustment             

  • 11:42 running ngc1039SU_IV OB to obtain data at SDT_Uspec,U-, I- and V-Bessel. Got AO corr from all filters.
  • 12:05 Slewing to Standard Star field SA95 and collimating --- it is getting light. I prepared an OB kuhn/TestOBs/SA95_UsUBVrg_rRIYziV Will try again with this one and other standards tonight (20100920).
  • Wait to take flats --- about 15-20 min before sunrise.
    • rI flats at 12:28. I has dark blob in lower right, between chips 1 and 2. See, lbcr123033 r looks OK. Count difference ~ 6220/5240 PA=0
    • At PA=180, take rI flats. lbcr123353. Still there.
    • At PA=180, go to SY filters. lbcr123629. Gone in Y filter.
    • Back to PA=180, rI sequence. lbcr123900 Still there on I filter (saturating also...)
    • Back to SY - gone.
    • Try VV.
    • Is there something on I-BESSEL filter. It is in Filter wheel #2 - this one is closest to M1. We went up to check. Steve is looking into L1 while I am rotating the filters. He sees a moth flying around inside. Took 1sec image at I-BESSEL - lbc.20100919.140751.fits with testcamera,but I cannot find this image. Pair of 0.5sec images through UI: 142503B and 142512 and 142540 at R. No sign of dark splotch on these images. It must have been the moth that Steve saw flying around behind L1 - I did not know it could get all the way to the filter - around L2 to L5, but it seems to have found a way.

  • Summary:
    • 52 stars for a pointing model; Check master/slave guiding; Checked for condensation on dewar window when long exposure (420 sec) taken with Red camera (see none and RH~65-70%).
    • One Red image shows 20 arcsec long trailed stars, but corresponding Blue image is good.
    • LBC Red Trackers RPC error.
    • Lost two red images taken in FPIA sequence - these did not appear in /newdata.
    • Want to check that offsets are as expected. (from data taken tonight)

  • Would have liked to check filter focus offsets, but seeing was not good tonight, at least in the 2nd half, after we had completed the collection of pointing model data.

Generating an LBC Pointing Model from the data taken on 20100919

The steps followed to generate a pointing model from these data taken on 20100919 are listed here:
  • Run LBTtools.LBC.qwcsauto on each of the LBC-Red pointing star images. qwcsauto matches a 5' x 5' region of the USNO-A2 catalog centered at the
coordinates of the Rotator Center to the stars detected on the image. When there are many stars, this works well, but when there are few stars, it took a few iterations, and as a last resort, I had to edit the cl script to allow it to match a larger region. For each star image, this output, in the working directory (/d1/kuhn/Pointing/20100919/), a fits file with a corrected WCS.
  • ls *.fits > 20100919ptfiles.list
  • mktpfile: Note that ctrlM = no. Otherwise, the lines are truncated midway through, the file is interpreted as a Gemini file, and the results using this are of order 10000 arcsec off.
imgtools> lpar mktpfile
        inlis = "pt20100919files.list" Input file list
    ptlogfile = "20100919.ptlog1" Pointing log file from PCS?
     (outfile = "20100919out")  Filename root for output data.
     (refXpix = 1080.)          Reference pixel X coordinate.
     (refYpix = 2912.)          Reference pixel Y coordinate.
       (ctrlM = no)             Trim the ^M at the end of each line?
     (verbose = yes)            Verbose output to screen?
      (ptlist = "")             
      (cplist = "")             
        (mode = "h")            
  • scp 20100919out"*" to telescope@lbtdu16. I created a directory, /home/telescope/okuhn/Pointing20100919/ for this.
  • Follow the commands listed on the screen output of mktpfile.
    • /lbt/tpt/bin/ptconv < 20100919out.ptconv.in > 20100919out.tpoint.in
    • paste 20100919out.tpoint.in 20100919out.aux_A1A.dat > 20100919out.tpoint.in2
    • /lbt/tpt/bin/tpoint and fit using IA IE NPAE CA AW AN TF TX
    • OUTMOD 20100919out.tpoint.out (The output file is 20100919out.tpoint.out - I have not yet copied this file to the mountain).
[telescope@lbtdu16 Pointing20100919]$ more 20100919out.tpoint.out
LBT Pointing Log File, 2010 Sep 19
T   52   4.8936    0.000   0.0000
  IA       -388.7827     8.54484
  IE        -35.7521     2.41471
  NPAE      +18.0420     9.80759
  CA        -58.7461    12.15142
  AW         -0.4340     0.79855
  AN        +20.4733     0.77708
  TF        +46.8804     5.94714
  TX        +10.2072     1.74946
END

  • File copied to /home/telescope/TCS/Configuration/PCS/LBCDX.20100919.ptmod and the two lines of PCSInstrument.conf
which referred to the LBC Pointing Model were edited to replace the name of the old one (LBCSX.20100521.ptmod) with this filename.

-- OlgaKuhn - 19 Sep 2010
  • Plot generated by TPOINT: fitting IA IE NPAE CA AW AN TF TX:
    20100919.tpoint1.jpg
Topic revision: r7 - 26 Sep 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback